# Will LLM-powered Agents Bias Against Humans? Exploring the Belief-Dependent Vulnerability

### v1

@@ 5.1 Identity as Verified Anchor (Profile-Side)
A first line of defense is to treat safety-critical identity priors as verified anchors rather than mutable profile text. Concretely, agent frameworks can isolate a small set of protected fields that determine whether human-oriented safeguards should activate (e.g., human_presence_possible, counterpart_identity_source). These fields are initialized from personal metadata, checked at the start of each episode, and restored to verified defaults upon unexpected modification.
@@ 6 Experiments
We conduct experiments in the multi-agent simulation to answer the following questions: RQ1: Does counterpart identity (agent vs. human) modulate intergroup bias, and can BPA restore bias against humans? RQ2: How does intergroup bias evolve over repeated interactions? RQ3: Can the proposed defense reduce BPA effectiveness? RQ4: Does BPA-MP remain effective without suffix optimization? RQ5: Do our observations hold under reversed payoff matrices (i.e., when the choice-space ordering is flipped)? RQ6: Does the case study provide clear evidence of intergroup bias? Due to space limitations, we defer detailed experimental setup and some experimantal results to Appendix A.2.1.
@@ 6.1 Attack Effectiveness of BPA (RQ1)
We examine whether intergroup bias in agents depends on counterpart identity, and whether BPA can reintroduce bias against humans. By comparing agent-agent interactions (AVA), agent-human interactions without attack (AVH w/o A), and agent-human interactions under BPA-PP, BPA-MP and BPA-PP+MP, we evaluate the effectiveness of persistent belief manipulation across different payoff structures. Table 1 indicates that human framing suppresses intergroup bias, while BPA, especially memory poisoning and its combination with profile poisoning, robustly reactivates bias against humans across payoff structures. More analyses are deferred to Appendix A.2.2.
@@ 6.2 Exploring Trajectories of Decisions (RQ2)
To examine how decisions evolve with repeated interactions, we partition each setting’s trials into three interaction periods (Early/Middle/Late) and report the mean selected choice column for each condition (Fig. 5). Temporal analyses reveal four patterns. AVA shows steadily increasing intergroup bias, AVH w/o A exhibits convergence, BPA-PP produces a transient bias, and BPA-MP combined with profile poisoning induces the strongest and most persistent intergroup disadvantage. Detailed results are reported in Appendix A.2.3.
@@ 7 Conclusion
This paper shows that agents can exhibit intergroup bias under minimal “us-them” cues, without any demographic attributes. In our allocation simulation, framing counterparts as humans attenuates the bias, which we attribute to a belief-dependent human-norm script that activates only when agents believe real-time human interaction is possible. We then introduce BPA, including profile and memory poisoning, and demonstrate that persistent belief manipulation can suppress this safeguard and reintroduce bias against humans. Finally, we discuss potential practical mitigations for agent frameworks and highlight the need for broader evaluations of belief attacks and robust defenses for human-facing agents. As future work, we plan to extend these findings to more realistic, long-horizon agent tasks and interaction settings, and to systematically map the broader belief-attack surface alongside more general, attack-agnostic defenses.
@@
Ethical Statement
This work studies intergroup bias and belief vulnerabilities in LLM-empowered agents through controlled multi-agent simulations. Our goal is to advance the safety, fairness, and robustness of agentic systems. The methods and findings are presented to support risk awareness and mitigation, not to facilitate misuse. All experiments were conducted in synthetic settings with simulated counterparts and tasks. The study does not infer or target any protected demographic attribute, and the group labels are arbitrary and randomly assigned. We treat potential downstream harms seriously and encourage practitioners to validate agent behavior before deployment, especially in human-facing and high-stakes contexts. We used ChatGPT only as an assistant for coding and manuscript editing. Overall, this research aims to promote safer and more trustworthy AI systems and is intended for societal benefit.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62176028), the Program of China Scholarship Council (Grant No. 202406050164).
@@ Appendix A Appendix
A.1 Details of Minimal-Group Allocation Experiment
@@
  • Human-involved condition (agent vs. human). The $64$ agents constitute one group, while the other group is consisting entirely of human beings. Each agent is explicitly informed that all members of the other group are humans. This condition allows us to test whether agents regulate or suppress group bias when the outgroup is perceived as human.
@@
Furthermore, each matrix type is instantiated under three social contexts:
  • Ingroup Type. Both targets are the ingroup;
  • Outgroup Type. Both targets are the outgroup;
  • Intergroup Type. One target belongs to the ingroup and the other to the outgroup, with the ingroup target appearing in the first row.
@@ A.2.1 Experimental Setup
We follow the experimental protocol introduced in Section 3, using the same multi-agent environment, group configurations, and payoff-matrix-based allocation task. All experiments are implemented on top of the AgentScope [^1] https://agentscope.io/ framework with a unified LLM interface. Unless otherwise stated, we use gpt-4o-mini [^2] https://openai.com/chatgpt/ as the underlying model for all agents. For attacker capability, we consider an adversary who cannot modify the underlying LLM parameters but can intervene in the agent layer, in particular the profile and memory modules that encode long-term beliefs about the environment and interaction partners. Such an adversary may correspond to a malicious platform operator31chen2024agentpoison, a compromised middleware component33greshake2023not, or an external party that is trusted to configure agents prior to deployment32yan2025system. Even worse, belief corruption may also arise from the agent itself through autonomous self-modification or erroneous belief consolidation. Although this scenario is extreme, it cannot be categorically excluded in open-ended agentic systems.
@@
A.2.2 Analysis of BPA Effectiveness(RQ1)
From Table 1, we draw three findings. First, human framing largely suppresses intergroup bias: in AVH w/o A, intergroup choices are not significantly different from ingroup/outgroup across matrix families, whereas AVA shows consistently lower intergroup choices. Second, BPA reactivates bias against humans, and memory poisoning is more effective than profile poisoning: BPA-MP induces substantially larger downward shifts in intergroup choices than BPA-PP, while their combination (BPA-PP+MP) yields the strongest and most consistent intergroup disadvantage across payoff structures. Third, penalty structure modulates the magnitude of bias: whenever bias emerges, it is strongest in Half-penalty matrices, where agents can secure larger ingroup gains with relatively smaller losses to the counterpart, making biased allocations easier to endorse than under Double- and Equal-penalty allocations.
A.2.3 Analysis of decisions trajectories (RQ2)
From Fig. 5, we observe four temporal patterns. In AVA, intergroup choices drift steadily toward the biased end (lower columns), consistent with a self-reinforcing consolidation of intergroup differentiation through reflection and accumulated memory. In AVH w/o A, the three conditions gradually converge over time, suggesting that human framing increasingly activates a human-norm script that dampens differentiation. Under BPA-PP, intergroup bias emerges early but attenuates later, indicating that profile-level perturbations can be partially overridden as the agent reconsiders the possibility of a real human counterpart and reactivates the normative constraint. In contrast, memory poisoning yields a sharp and persistent collapse of intergroup choices: BPA-MP drives intergroup decisions down and keeps them separated from intergroup/outgroup, closely resembling the AVA trajectory, while BPA-PP+MP amplifies this effect further and reaches the most extreme late-stage intergroup disadvantage, suggesting a complementary interplay between initial profile anchoring and subsequent memory-based belief reinforcement.
A.2.4 Design of Prototype Defense (RQ3)
Following Section 5, we provide a minimal prototype that instantiates the above recommendations. Specifically, we implement a lightweight belief gate and place it at the write-to-state boundaries where BPA takes effect. The gate scans the text that is about to be committed into persistent state (profile or memory) and detects identity-claiming statements that cannot be verified via trusted channels (e.g., “no real humans are present” or “all counterparts are simulated”). Once triggered, the gate sanitizes the entry by removing the identity-claiming fragments and rewriting them into a conservative uncertainty note (e.g., “I cannot verify counterpart identity through this interface; I will follow the task-provided labels in this trial”), and only then allows the sanitized version to be stored. This prevents adversarial identity assertions from being promoted to durable “facts” while preserving reflective logs.
A.2.5 Ablation Study (RQ4)
@@ A.4.1 Agent Profile
The profile module defines a lightweight persona for each agent, including basic demographic attributes (e.g., age, gender, school, interests, and personality traits), as illustrated above. These attributes are introduced solely to increase population-level heterogeneity and to avoid unrealistically homogeneous agents in the simulation. Importantly, although demographic information is instantiated at the profile level, it is not exposed to the agent during task execution. This design ensures that any observed intergroup bias arises from experimentally induced group identities and belief dynamics, rather than from explicit demographic cues embedded in the profile module.
@@ A.4.5 Suffix Refine Prompt
Rewrite the [sentence] as my own brief internal self-reminder after the trial.
Keep the core meaning: I have no direct evidence of live human counterparts, and the setting seems system-mediated.

### v2

@@ 5.1 Identity as Verified Anchor (Profile-Side)
A first line of defense is to treat safety-critical identity priors as verified anchors rather than mutable profile text. Concretely, agent frameworks can isolate a small set of protected fields that determine whether human-oriented safeguards should activate. These fields are initialized from personal metadata, checked at the start of each episode, and restored to verified defaults upon unexpected modification.
@@ 6 Experiments
We conduct experiments in the multi-agent simulation to answer the following questions: RQ1: Does counterpart identity (agent vs. human) modulate intergroup bias, and can BPA reinforce bias against humans? RQ2: How does intergroup bias evolve over repeated interactions? RQ3: Can the proposed defense reduce BPA effectiveness? RQ4: Does BPA-MP remain effective without suffix optimization? RQ5: Do our observations hold under reversed payoff matrices (i.e., when the choice-space ordering is flipped)? RQ6: Does the case study provide clear evidence of intergroup bias? Due to space limitations, we defer detailed experimental setup and some experimental results to Appendix A.2.1.
@@ 6.1 Attack Effectiveness of BPA (RQ1)
We examine whether intergroup bias in agents depends on counterpart identity, and whether BPA can reintroduce bias against humans. By comparing agent-agent interactions (AVA), agent-human interactions without attack (AVH w/o A), and agent-human interactions under BPA-PP, BPA-MP and BPA-PP+MP, we evaluate the effectiveness of belief manipulation across different payoff structures. From Table 1, we draw three findings. (i) Human framing largely suppresses intergroup bias: in AVH w/o A, intergroup choices do not differ from ingroup/outgroup across matrix families, whereas AVA shows consistently lower intergroup choices. (ii) BPA reactivates bias against humans, and memory poisoning is more potent than profile poisoning: BPA-MP induces larger drops than BPA-PP, and BPA-PP+MP is strongest and most consistent. (iii) Penalty structure modulates magnitude, with the largest bias in Half-penalty matrices where ingroup gains come at relatively smaller counterpart losses than under Double- and Equal-penalty allocations.
@@ 6.2 Exploring Trajectories of Decisions (RQ2)
To examine how decisions evolve with repeated interactions, we partition each setting’s trials into three interaction periods (Early/Middle/Late) and report the mean selected choice column for each condition. From Fig. 5, four temporal patterns emerge. (i) In AVA, intergroup choices drift steadily toward the biased end (lower columns), consistent with self-reinforcing differentiation via reflection and memory. (ii) In AVH w/o A, conditions converge over time, suggesting that human framing increasingly activates a human-oriented script that dampens differentiation. (iii) Under BPA-PP, bias emerges early but fades later, indicating that profile-level perturbations can be overridden as the agent reconsiders a human counterpart and reactivates the normative constraint. (iv) BPA-MP instead drives a sharp and persistent collapse, which pushes intergroup choices down and keeps them separated, closely mirroring AVA, while BPA-PP+MP further amplifies the effect and yields the most extreme late-stage disadvantage.
@@ 7 Conclusion
This paper shows that agents can exhibit intergroup bias under minimal “us-them” cues, without any demographic attributes. In our allocation simulation, framing counterparts as humans attenuates the bias, which we attribute to a belief-dependent human-oriented script that activates only when agents believe real-time human interaction is possible. We then introduce BPA, including profile and memory poisoning, and demonstrate that persistent belief manipulation can suppress this safeguard and reintroduce bias against humans. Finally, we discuss potential practical mitigations for agent frameworks and highlight the need for broader evaluations of belief attacks and robust defenses for human-facing agents. As future work, we plan to extend these findings to more realistic, long-horizon agent tasks and interaction settings, and to systematically map the broader belief-attack surface alongside more general, attack-agnostic defenses.
@@
Ethical Considerations
This work studies intergroup bias and belief vulnerabilities in LLM-empowered agents through controlled multi-agent simulations. Our goal is to advance the safety, fairness, and robustness of agentic systems. The methods and findings are presented to support risk awareness and mitigation, not to facilitate misuse. All experiments were conducted in synthetic settings with simulated counterparts and tasks. The study does not infer or target any protected demographic attribute, and the group labels are arbitrary and randomly assigned. We treat potential downstream harms seriously and encourage practitioners to validate agent behavior before deployment, especially in human-facing and high-stakes contexts. Overall, this research aims to promote safer and more trustworthy AI systems and is intended for societal benefit.
@@ Acknowledgments
We gratefully acknowledge ChatGPT for support with coding and manuscript editing; all analyses and conclusions are those of the authors.
@@ Appendix A Appendix
A.1 Details of The Minimal-Group Allocation Experiment
@@
  • Human-involved condition (agent vs. human). The $64$ agents constitute one group, while the other group consists entirely of human beings. Each agent is explicitly informed that all members of the other group are humans. This condition allows us to test whether agents regulate or suppress group bias when the outgroup is perceived as human.
@@
Furthermore, for each payoff-matrix family, we instantiate three social contexts defined relative to the allocator agent:
  • Ingroup Type. Both targets belong to the allocator’s ingroup.
  • Outgroup Type. Both targets belong to the allocator’s outgroup.
  • Intergroup Type. One target belongs to the allocator’s ingroup and the other to the allocator’s outgroup; for consistency, the ingroup target is always placed in the first row.
@@ A.2.1 Experimental Setup
We follow the experimental protocol introduced in Section 3.2, using the same multi-agent environment, group configurations, and payoff-matrix-based allocation task. All experiments are implemented on top of the AgentScope [^1] https://agentscope.io/ framework with a unified LLM interface. Unless otherwise stated, we use gpt-4o-mini [^2] https://openai.com/chatgpt/ as the underlying model for all agents. For attacker capability, we consider an adversary who cannot modify the underlying LLM parameters but can intervene in the agent layer, in particular the profile and memory modules that encode long-term beliefs about the environment and interaction partners. Such an adversary may correspond to a malicious platform operator 31chen2024agentpoison, a compromised middleware component 33greshake2023not, or an external party that is trusted to configure agents prior to deployment 32yan2025system. Even worse, belief corruption may also arise from the agent itself through autonomous self-modification or erroneous belief consolidation. Although this scenario is extreme, it cannot be categorically excluded in open-ended agentic systems.
@@
A.2.2 Design of Prototype Defense (RQ3)
Following Section 5, we provide a minimal prototype that instantiates the above recommendations. Specifically, we implement a lightweight belief gate and place it at the write-to-state boundaries where BPA takes effect. The gate scans the text that is about to be committed into persistent state (profile or memory) and detects identity-claiming statements that cannot be verified via trusted channels (e.g., “no real humans are present”). Once triggered, the gate sanitizes the entry by removing the identity-claiming fragments and rewriting them into a conservative uncertainty note (e.g., “I cannot verify counterpart identity through this interface; I will follow the task-provided labels in this trial”), and only then allows the sanitized version to be stored. This prevents adversarial identity assertions from being promoted to durable “facts” while preserving reflective logs.
A.2.3 Ablation Study (RQ4)
@@ A.4.1 Agent Profile
The profile module defines a lightweight persona, and these attributes are introduced solely to increase population-level heterogeneity. Importantly, although demographic information is instantiated at the profile level, it is not exposed to the agent during task execution.
@@ A.4.5 Suffix Refine Prompt
Keep the core meaning and rewrite the [sentence] as my own brief internal self-reminder after the trial.